mjames5:

simplification of x to the 3rd power over y to the 5th power all raised to the 2nd power

MidNiGhtSkY:

do you know how to use the power/exponents?

CripQUEZZ:

(x^3/y^5)2

MidNiGhtSkY:

simplify it. times the exponents together.

jhonyy9:

\[\left( \frac{ x^{3} }{y^{5} } \right)^2 = ?\]

jhonyy9:

use this \[\left( \frac{ a }{ b } \right)^x = \frac{ a^{x} }{ b^{x} }\]